OpenStudy (anonymous):

When planning his crops, Farmer Sue knows that her 15 acres can support apples and pecans. She wants to make $8,500 from her crops. She can make $1,050 per acre of apples (the variable a) and $2,500 per acre of pecans (the variable p). Which equation below would be a constraint in her system of equations?

OpenStudy (anonymous):

1,050a + 2,500p = 8,500 a + p = 15 a - p = 15 2,500p - 1,050a = 8,500

OpenStudy (anonymous):

I know it is either a or b but, in my opinion they are both true so what would be the constraint